2) A train travels 300 km in the same time that a plane covers 800 km. If the speed of the plane is 20 km per hr less than 3 times the speed of the train, find both speeds.
help me please!

Answer by checkley77(12844)

D=RT
300=RT FOR THE TRAIN
OR: T=300/R
800=(3R-20)T FOR THE PLANE
OR: T=800/(3R-20)
BECAUSE THE TIMES ARE EQUAL WE HAVE:
300/R=800/(3R-20)
800R=300(3R-20)
800R=900R-6,000
800R-900R=-6,000
-100R=-6,000
R=-6,000/-100
R=60 KMH FOR THE TRAIN.
3*60-20=180-20=160KMH FOR THE PLANE.
PROOF:
300/60=800/160
5=5 HOURS FOR THE TRIPS.